OpenStudy (anonymous):

Find the area of the circle if its circumference is 30m. (Please show the solution. I need the solution. Thanks)

OpenStudy (diyadiya):

Circumference = \(2 \times \pi \times radius \) 30 = \(2 \times 3.14 \times radius \) can you find radius now?

OpenStudy (diyadiya):

After finding radius, find the area of the circle using the formula Area of a circle = \( \pi \times (radius)^2 \)

OpenStudy (diyadiya):

is this clear?

OpenStudy (diyadiya):

Sure? Did you get the answer? :D You're Welcome!